TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Create new blog entry for Point 8


Created by Main.MircoCoccoli on 2006-12-20 - 12:01

RAT Preparation Cool Down sect 78 - Dec 20th

PDF RAT_20061220.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x20x11x10/RAT_20061220.pdf

Next RAT meeting Tuesday, January 9th 8:30 @ P8 2889-R-009

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 15.12.06 DFBXG non-conformity on the current lead heater cables (P.Chambouvet-J-C.Guillaume)
Cryogenics
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
DFB Commissioning
  • 29.11.06 DFBX Pumping of Current Leads insulation vacuum (S.Feher)
13 kA EE Systems
  • 16.12.06 Calibration of the Arc Detectors on the 13 kA EE System (K.Dahlerup-Petersen)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

MircoCoccoli - 20 Dec 2006 - 11:11

Edit, attach file or add comment


Created by Main.RobertoSaban on 2006-12-20 - 09:52

Incident Summary Report - Short on the MB circuit during TP4B

During the Electrical Quality Assurance (TP4B) of sector 78 a short to ground was identified. The short chronological report of the events is attached.

PDF short_21R7.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x20x09x48/short_21R7.pdf

PowerPoint FSavary.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x20x09x48/FSavary.ppt
Report to MARIC of the repair sequence
PowerPoint DBozzini.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x20x09x48/DBozzini.ppt
Report to MARIC of the ELQA tests

Comments:

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-19 - 12:00

Preparation Cool-Down Sector 78-RAT 19th Dec

PDF RAT_20061219.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x19x11x13/RAT_20061219.pdf

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 15.12.06 DFBXG non-conformity on the current lead heater cables (P.Chambouvet-J-C.Guillaume)
Cryogenics
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
DFB Commissioning
  • 29.11.06 DFBX Pumping of Current Leads insulation vacuum (S.Feher)
13 kA EE Systems
  • 16.12.06 Calibration of the Arc Detectors on the 13 kA EE System (K.Dahlerup-Petersen)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

MircoCoccoli - 19 Dec 2006 - 11:56

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-12-18 - 12:28

Preparation Cool-Down Sector 78-RAT 18thDec

PDF RAT_20061218.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x18x12x27/RAT_20061218.pdf

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-12-15 - 12:15

Preparation Cool-Down Sector 78-RAT 15th Nov

PDF RAT_20061215.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x15x12x14/RAT_20061215.pdf

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-12-14 - 18:25

Preparation Cool-Down Sector 78-RAT 14th Nov

PDF RAT_20061214.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x14x18x20/RAT_20061214.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2006-12-14 - 15:15

60A converter tests - Thu. 14th Dec.

Advancement of today:

  • connection of the half arc L8 finished: ready for ELQA.doc
  • TS-EL, resolution of the problem on the short circuit insulation in R7, no problems for co activities.

Tomorrow activities:

  • TS-EL, still on resolution of the problem on the short circuit insulation in R7, no problems for co activities.
  • If it is possible: El_QA.doc 120A circuits in RR77 and UA83, no problems for co activities.

Edit, attach file or add comment


Created by Main.BorisBellesia on 2006-12-13 - 20:28

60A converter tests - Wed. 13th Dec.

JPEG P1000122.JPG:
P1000122.JPG

Advancement of today:

  • Pre-connection of the cable of the current leads: on C28-C32_L8
  • All 60A cables disconnected PC side:
  • Problem on the short circuit insulation on the 60A cables in R7 (see pic)

Tomorrow activities:

  • TS-EL, resolution of the problem on the short circuit insulation in R7, no problems for coactivities.
  • if it is possible: El_QA.doc 120A circuits in RR77 and UA83, no problems for coactivities.

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-13 - 17:25

RAT Preparation Cool Down sec78 - Dec 13

PDF RAT_20061213.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x13x17x22/RAT_20061213.pdf

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 12.12.06 Short-to-ground in QBBI.A21R7 (Lyra side) (D.Bozzini, F.Seyvet)
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBXG (P.Chambouvet)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Vacuum
  • 12.12.06 One Pressure Gauge in DFBs cannot be read (F.Millet)
Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO (F.Millet)
  • 22.11.06 Control of quench valve (F.Millet)
DFB Commissioning
  • 12.12.06 Repositioning of 1 DC cable on DFBAO and of 2 DV cables on DFBAN (all 13kA) (J.C.Guillaume)
  • 08.12.06 DFBXG 100V test on current leads (S.Feher)
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk50 (J-C.Guillaume & R.Jones)

MircoCoccoli - 13 Dec 2006 - 17:22

Edit, attach file or add comment


Created by Main.BorisBellesia on 2006-12-12 - 20:28

60A converter tests - Tuesday 12th Dec.

Advancement Tue. 12 Dec.:

  • AB-PO activities ended in the whole arc
  • Polarity tests ended in the whole arc

Activities Wed. 13th Dec.:

  • TS-EL pre-conneciton in C32-34_L8 / C 12-34_R7; no problems for coactivities
  • if it is possible: El_QA.doc in C18-28 _L8; co activities limited - a trolley is used

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-12 - 12:40

Preparation for Cool Down RAT sec78 - Dec 12

PDF RAT_20061212.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x12x12x38/RAT_20061212.pdf

Comments:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBXG (P.Chambouvet)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Vacuum
  • 12.12.06 One Pressure Gauge in DFBs cannot be read (F.Millet)
Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO (F.Millet)
  • 22.11.06 Control of quench valve (F.Millet)
DFB Commissioning
  • 08.12.06 DFBXG 100V test on current leads (S.Feher)
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk50 (J-C.Guillaume & R.Jones)

MircoCoccoli - 12 Dec 2006 - 12:38

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-11 - 16:08

Preparation for Cool Down RAT - Dec 11

PDF RAT_20061211.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x11x16x04/RAT_20061211.pdf

Comments:

AC non-conformity

  • 07.12.06 Non conformity in DFBMH, DFBAN and DFBXG current lead heaters (P.Chambouvet-J-C.Guillaume)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBXG (P.Chambouvet)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO (F.Millet)
  • 22.11.06 Control of quench valve (F.Millet)
DFB Commissioning
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk50 (J-C.Guillaume & R.Jones)

Main.MircoCoccoli - 11 Dec 2006 - 16:07

Edit, attach file or add comment


Created by Main.BorisBellesia on 2006-12-08 - 17:54

60A converter tests - Friday 7th Dec.

Advancement of today:

  • AB-PO activities on cells C24-30_R7
  • 4 hr run in C12_R7-C22_R7 and C32-C34_R7 and C28-C32_L8 simultaneously (shortened version of the 24 hr run)
  • AB-PO main activity ended today;
  • Polarity tests on cells C24-26_R7

Monday activities:

  • AB-PO - tests on faulty cells: - C34_L8 - 1 converter in C32_R7 - 1 converter in C12_R7 - no problems for coactivities.
  • Polarity and conneciton in C28-32_L8 / C30-34_R7; no problems for coactivities
  • ELQA in C18-26_L8; co activities limited - a trolley is used

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-08 - 10:27

Preparation for Cool Down RAT - Dec 8

PDF RAT_20061208.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x08x10x17/RAT_20061208.pdf

IMPORTANT: PLEASE COMMENT ON THE OPEN ISSUES IF YOU FIND THEM INCORRECT, INCOMPLETE OR NOT-PERFECTLY UPDATED.

AC non-conformity

  • 07.12.06 Non conformity in DFBMH, DFBAN and DFBXG current lead heaters (P.Chambouvet-J-C.Guillaume)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling

  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBXG (P.Chambouvet)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Vacuum -

Cryogenics

  • 22.11.06 Cryo-valves remote control in CCC by AB/CO (F.Millet)
  • 22.11.06 Control of quench valve (F.Millet)
DFB Commissioning
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk50 (J-C.Guillaume & R.Jones)

MircoCoccoli - 08 Dec 2006 - 10:17

Edit, attach file or add comment


Created by Main.BorisBellesia on 2006-12-07 - 19:02

60A converter tests - Thursday 7th Dec.

Comments:

Advancement of today:

  • AB-PO activities on cells C12_R7 - C16-22_R7
  • Polarity tests on 120A converter alimenting cells C6-7-8-9-10_R7
  • Disconnection of the circuits of the 120A converter alimenting cells C6-7-8-9-10_R7
  • MME activities postponed to Monday

Tomorrow activities:

  • AB-PO activities on cells C24-30_R7
  • 4 hr run in C12_R7-C22_R7 and C32-C34_R7 and C28-C32_L8 simultaneously (shortened version of the 24 hr run)
  • ELQA tests - whole day / no co-activities

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-07 - 12:37

Preparation for Cool Down sec78 - Dec 7

PDF RAT_20061207.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x07x12x29/RAT_20061207.pdf

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 30.11.06 Non conformity: 4 temperature sensor cables need repair (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC (Flohe)
Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
DFB Commissioning
  • 04.12.06 Re-positioning collars used to fix the WRL to the CLs (all DFBs) (A.Perin-D.Bozzini-F.Millet)
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: next(w48): DFBAO HC module, DFBMA, DFBMC, DFBX))
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1. To be discussed on next RAT.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)

MircoCoccoli - 07 Dec 2006 - 12:36

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-06 - 15:17

Preparation for Cool Down sec78 - Dec 6

PDF RAT_20061206_wslides_wList.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x06x15x14/RAT_20061206_wslides_wList.pdf

Comments:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 30.11.06 Non conformity: 6 temperature sensor cables need repair (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC (Flohe)
Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
DFB Commissioning
  • 04.12.06 Re-positioning collars used to fix the WRL to the CLs (all DFBs) (A.Perin-D.Bozzini-F.Millet)
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: next(w48): DFBAO HC module, DFBMA, DFBMC, DFBX))
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1. To be discussed on next RAT.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)

Main.MircoCoccoli - 06 Dec 2006 - 15:16

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-05 - 13:50

Preparation for Cool Down sec78 - 5th Dec 06

PDF RAT_20061205_wslides.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x05x13x47/RAT_20061205_wslides.pdf
Minutes with Presentation by D.Bozzini attached

Open Issues:

AC non-conformity

  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 3 cables are in repair-Wed.29th)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 04.12.06 Access & Safety Cable Installation (full sector) (J-C.Guillaume; end by Dec 7th)
  • 30.11.06 Non conformity: 6 temperature sensor cables need repair (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC, DFBAN (Flohe)
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 04.12.06 Re-positioning collars used to fix the WRL to the CLs (all DFBs) (A.Perin-D.Bozzini-F.Millet)
  • 29.11.06 Pumping of insulation vacuum for DFBX
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

-- Main.MircoCoccoli - 05 Dec 2006 - 13:47

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-04 - 19:03

Preparation for Cool Down sec78 - RAT 4th Dec

PDF RAT_20061204.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x04x17x39/RAT_20061204.pdf

PDF Slide_w49.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x04x17x39/Slide_w49.pdf
Week 49 Usage Table

Open Issues:

AC non-conformity

  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 3 cables are in repair-Wed.29th)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 04.12.06 Access & Safety Cable Installation (full sector) (J-C.Guillaume; end by Dec 7th)
  • 30.11.06 Non conformity: 6 temperature sensor cables need repair (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC, DFBAN (Flohe)
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 04.12.06 Re-positioning collars used to fix the WRL to the CLs (all DFBs) (A.Perin-D.Bozzini-F.Millet)
  • 29.11.06 Pumping of insulation vacuum for DFBX
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: next(w48): DFBAO HC module, DFBMA, DFBMC, DFBX))
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1. To be discussed on next RAT.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

-- Main.MircoCoccoli - 04 Dec 2006 - 17:40

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-03 - 23:20

Sector 78 - Week 49 - Zones Usage

This tentative planning will be discussed at Monday 4th RAT meeting in Point 8.

Thanks, Mirco

PDF Plan78w49.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x03x23x13/Plan78w49.pdf

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-01 - 13:16

RAT - Preparation for Cool Down sec 78 - 1 Dec 06

PDF RAT_20061201.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x01x13x13/RAT_20061201.pdf

Open Issues:

AC non-conformity

  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 3 cables are in repair-Wed.29th)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 30.11.06 Non conformity: 6 temperature sensor cables need repair (A. Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC, DFBAN (Flohe)
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 20.11.06 DFB instrumentation cables to be connected (A.Suraci: 80% done)
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 29.11.06 Definition of owner of activity and pumping needs for current leads insulation vacuum (DFBX)
  • 29.11.06 Pumping of insulation vacuum for DFBX
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: next(w48): DFBAO HC module, DFBMA, DFBMC, DFBX))
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1. To be discussed on next RAT.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

-- Main.MircoCoccoli - 01 Dec 2006 - 13:13

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2006-12-11 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login